OpenStudy (angelberina):

A Florida citrus grower estimates that if 60 orange trees are planted, the average yield per tree will be 400 oranges. The average yield per tree will decrease by 4 oranges for each additional tree planted on the same acreage. How many trees should the grower plant to maximize the total yield?

OpenStudy (dumbcow):

Total yield = Avg yield * num of Trees Avg yield is 400 when n=60 but goes down by 4 for each additional tree ---> 400 - 4(n-60) \[Y = n(400-4(n-60))\] \[Y = 640n -4n^2\] Maximize by setting derivative equal to 0 \[\frac{dY}{dn} = 640 - 8n = 0\] \[n = 80\]